Question 1 The weight of a package delivered by a courier company is normally distributed with mean 15 kg and standard deviation 1.75 kg.

(a) There are 9% of the packages weigh heavier than w kg. Find the value of w. (4 marks)


(b) Suppose the shipping fee set by the courier company is $150 basic charge, together with an additional charge of $16 per kg of the package. Find the following summary statistics of the shipping fee: (I) Mean (II) Standard deviation (III) Variance (IV) Median (V) 91st percentile (6 marks)


1
Expert's answer
2021-12-22T09:45:11-0500

a)

"z=\frac{x-\mu}{\sigma}"


"P(x>w)=1-P(x<w)=0.09"

"P(x<w)=1-0.09=0.91"

"z=1.34"

"P(x<w)=P(z<1.34)=\frac{w-15}{1.75}=0.91"


"w=0.91\cdot1.75+15=16.6" kg


b)

X is cost of package

i)

mean:

"\mu_1=150+16\mu=150+16\cdot15=\$390"


iii)

Variance of weight:

"\sigma^2=\frac{\sum (x_i-\mu)^2}{n}=1.75^2"

Variance of fee:

"\sigma_1^2=\frac{\sum 16^2(x_i-\mu)^2}{n}=16^2\cdot1.75^2=784"


ii)

Standard deviation of fee:

"\sigma_1=\sqrt{784}=28"


iv)

The median is the middle point in a dataset—half of the data points are smaller than the median and half of the data points are larger.

for normal distribution:

median = mean = 389


v)

91st percentile is a score below which a 91% of scores in its frequency distribution falls

for normal distribution:

for "P=0.91" :

"z=1.34"

"P(z<1.34)=0.91"


"z=\frac{X-\mu_1}{\sigma_1}=\frac{X-390}{28}=1.34"


91st percentile of fee:

"X_{91}=1.34\cdot28+390=427.52"